An airplane leaves los angeles for Denver at a speed of 450 mph.Thirty min later,a plane going from Denver yo los Angeles leaves Denver,which is 850 miles from los Angeles,at a speed of 520 mph.When they meet, how far are they from Denver?
Found 2 solutions by mananth, graphmatics:
Answer by mananth(16949)

Let them meet at distance x from Denver
LA to Denver - speed = 450 mph
Denver to LA = 520 mph ( 30 minutes later)
Total distance = 850 miles
d=rt
When the second plane starts the first plane has already traveled 450*1/2 = 225 miles
so the distance between them will be 625 miles
add up their speeds = 520+450 = 970 mph
time = d/r
time = 625/970
= 0.64 hours
Distance from denver = 0.64*520= 334.8 miles
m.ananth@hotmail.ca


Answer by graphmatics(170)

After 30 minutes the first plane has traveled 225 miles from Los Angeles so there are 850-225=625 miles left to Denver. Let t be the time in hours before the 2 planes meet. The first plane will travel 450*t miles. The second plane will travel 520*t miles. The sum of the 2 distances must be 625 miles so. 450*t+520*t=625. 970*t=625. t=625/970. t=0.644 hours. The distance from Denver is 520*0.644=334.88 miles.